# Physics-Guided Multi-Fidelity DeepONet for Data-Efficient Flow Field Prediction

### v1

@@ 7 Conclusion
This study presents a novel multi-fidelity DeepONet (MF-DeepONet) framework for efficient spatio-temporal flow field prediction with reduced high-fidelity data requirements. Through strategic architectural enhancements and physics-guided sampling, we address critical challenges in operator learning for fluid dynamics applications. Our merge network architecture enables more complex feature interactions between operator and coordinate spaces, achieving a 50.4% reduction in prediction error compared to traditional dot-product approaches. The implementation of temporal positional encoding and efficient point-based sampling strategies further improves prediction accuracy by 7.57% while dramatically reducing training time by 96%. The transfer learning-based multi-fidelity framework leverages knowledge from pre-trained low-fidelity models to guide high-fidelity predictions, demonstrating a 43.7% improvement in accuracy compared to single-fidelity training. Most significantly, our time-derivative guided sampling strategy demonstrates two key advantages: (1) it reduces prediction error by 20.7% compared to conventional uniform sampling when using the same amount of high-fidelity data, and (2) it requires only 60% of the high-fidelity data to achieve comparable accuracy to models trained with full data using uniform sampling—a crucial advancement for computationally expensive simulations.
@@
Despite these promising results, our framework has several limitations that warrant further investigation. Although highly effective for our flow problems, the time-derivative guided sampling strategy requires broader validation across diverse physical systems with different temporal characteristics. Additionally, our current transfer learning approach—which completely freezes branch and trunk networks—may be overly restrictive; more flexible layer-freezing strategies based on data similarity metrics could potentially yield further improvements. Furthermore, the computational overhead introduced by automatic differentiation in temporal derivative calculations presents opportunities for developing more efficient sampling implementations. Lastly, to enhance practical applicability in industrial settings, extending the framework to accommodate irregular domains, complex geometries, and moving boundaries represents an important direction for future work.
@@
Future research should focus on developing adaptive physics-guided sampling criteria beyond temporal derivatives to improve cross-domain generalizability. An automated layer-wise similarity analysis between low-fidelity and high-fidelity features could enable more nuanced transfer learning approaches with selective network component freezing. Computational efficiency could be enhanced through surrogate-based derivative estimation or alternative physics-guided sampling metrics that avoid expensive automatic differentiation calculations. Finally, incorporating geometric information directly into the sampling strategy would enable effective point selection in complex domains, broadening the framework’s applicability across diverse scientific disciplines where high-fidelity data acquisition remains challenging. These advancements would further establish MF-DeepONet as a practical solution for real-time flow prediction in data-constrained environments.
@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (2018R1A5A7025409), and the Ministry of Science and ICT of Korea grant (No. 2022-0-00969, No. 2022-0-00986, and No. RS-2024-00355857), the Ministry of Trade, Industry & Energy grant (RS-2024-00410810), and the National Research Council of Science & Technology (NST) grant by the Korea government (MSIT) (No. GTL24031-000)

### v2

@@ 8 Conclusion
This study presents a novel multi-fidelity DeepONet (MF-DeepONet) framework for efficient spatio-temporal flow field prediction with reduced high-fidelity data requirements. Through strategic architectural enhancements and physics-guided subsampling, we address critical challenges in operator learning for fluid dynamics applications. Our merge network architecture enables more complex feature interactions between operator and spatio-temporal coordinate spaces, achieving a 50.4% reduction in prediction error compared to traditional dot-product approaches. The implementation of temporal positional encoding and efficient point-based sampling strategies further improves prediction accuracy by 7.57% while dramatically reducing training time by 96%. The transfer learning-based multi-fidelity framework leverages knowledge from pre-trained low-fidelity models to guide high-fidelity predictions, demonstrating a 43.7% improvement in accuracy compared to single-fidelity training. Most significantly, our time-derivative guided subsampling strategy demonstrates two key advantages: (1) it reduces prediction error by 20.7% compared to conventional uniform sampling when using the same amount of high-fidelity data, and (2) it achieves prediction accuracy comparable to a model trained on the complete set of available high-fidelity data, while using only 60% of those training data—a crucial advancement for computationally expensive simulations. The robustness of these contributions is further validated on a second, more complex dataset with higher spatial resolutions and more intricate flow physics, where our proposed MF-DeepONet framework and time-derivative subsampling strategy continued to demonstrate superior or competitive performance against conventional benchmarks.
@@
Despite these promising results, our framework also has limitations that warrant further investigation. First, the current transfer learning approach—which completely freezes branch and trunk networks—may be overly restrictive; more flexible layer-freezing strategies based on data similarity metrics could potentially yield further improvements. Additionally, while the time-derivative guided subsampling strategy proved effective for our flow problems, it requires broader validation across diverse physical systems with different temporal characteristics. Furthermore, its reliance on automatic differentiation introduces computational overhead, presenting opportunities for developing more efficient sampling implementations. Lastly, for broader industrial applicability, extending the framework to accommodate irregular domains, complex geometries, and moving boundaries represents an important direction for future work.
@@
Future research will focus on several key areas to build upon the findings of this study. First, to improve the multi-fidelity framework itself, an automated layer-wise similarity analysis between low-fidelity and high-fidelity features could enable more nuanced transfer learning approaches with selective network component freezing. The observed training instability in some cases—attributable to the initial training of LF-DeepONet without guidance from HF data—highlights the need for refining the training strategy, which could serve as a promising direction for future research. Another important extension would be to investigate the robustness of MF-DeepONet when the LF and HF datasets are less strongly correlated, such as cases involving different Reynolds numbers or slight geometric domain changes. This would clarify the level of physical similarity required for effective transfer and broaden the framework’s applicability in real-world scenarios. Furthermore, to enhance predictive robustness for time-evolving problems, future work will investigate the incorporation of physics-informed loss functions that more explicitly enforce the underlying Navier-Stokes dynamics during training. Second, to advance the sampling strategy, research can focus on developing physics-guided subsampling criteria beyond temporal derivatives—such as incorporating geometric information—to improve cross-domain generalizability and enable effective point selection in complex domains. The computational efficiency of these methods could also be enhanced through surrogate-based derivative estimation or alternative physics-guided metrics that avoid expensive automatic differentiation calculations. These advancements would further establish MF-DeepONet as a practical solution for real-time flow prediction in data-constrained environments.
@@ Acknowledgments
This work was supported by the Ministry of Science and ICT of Korea grant (No. RS-2024-00355857), and the National Research Council of Science & Technology (NST) grant by the Korea government (MSIT) (No. GTL24031-000)